OpenStudy (anonymous):

30 + 30x = 25 + 55x solve for x can someone help?

OpenStudy (ivettef365):

first combine like terms, so when you switch from one side to another just change the sign as follows: 30 -25 = 55x - 30x can you take it from here ?

OpenStudy (anonymous):

I did that and I got 1/5 but it didn't fit the scenario

OpenStudy (ivettef365):

I know why, you did the equation wrong, it should be 30x +30 = 25x + 55 try it now

OpenStudy (ivettef365):

so now you have 30x -25x = 55 - 30 what do you get ?

OpenStudy (ivettef365):

@RICARDOismyfirstname ???

OpenStudy (anonymous):

im working on it right now

OpenStudy (anonymous):

okay so from there this is what i got: 30x + 30 = 25x + 55 30x - 25x = 55 - 30 5x = 25 -- -- 5 5 x = 5 ============== Then put it back into the equation 30(5) + 30 = 25(5) + 55 150 + 30 = 125 + 55 180 = 180

OpenStudy (phi):

your only mistake was when you set up the equations the first time. 25 + 55x would mean 55 per yard plus 25 you really want 25 x +55 (meaning 25 per yard) plus 55

OpenStudy (anonymous):

yeah @phi i got it mixed up when i was thinking of it, im pretty sure I got it now! thank you guys!!!!

OpenStudy (ivettef365):

that is correct